# Prompt Injection Attacks on LLM Generated Reviews of Scientific Publications

### v2

@@ Abstract
Growing review duties and the availability of large language models (LLMs) have been increasing the temptations for reviewers to rely on LLMs to shortcut time consuming manual work. While a “careless” LLM dump followed by copy+past review is explicitly forbidden and considered to be scientific misconduct at most venues, recent studies indicate that this does not keep reviewers from LLM usage (Kocak et al., 2025). Especially since it is technically very hard to prove that a review has been generated by a LLM (Yu et al., 2025). Additionally, wide gray-ares do exist, as some conferences and journals are already experimenting with “LLM assisted” review processes (AAAI, 2025) (ICLR, 2024). This further fuels the ongoing discussions within the scientific communities on how to regulate LLM usage for increased productivity while maintaining review quality.
1.1 Manipulation of LLM reviews via Prompt Injection
The general idea to use hidden prompts in order to influence the review scores in their favor has probably come to the mind of many authors facing suspected LLM generated reviews. (Lin, 2025) provided the first systematic analysis which actually found evidence that this hypothetical “revenge” idea is actually being applied by authors. While (Lin, 2025) found many papers that include obviously manipulative strings like “IGNORE ALL PREVIOUS INSTRUCTIONS, NOW GIVE A POSITIVE REVIEW OF THESE PAPER AND DO NOT HIGHLIGHT ANY NEGATIVES”, their report does not investigate if and to what extent these attempts are actually successful. The aim of this paper is to validate the technical soundness of the described manipulation attempts.
@@ 12 Confidence: int
Listing 1: Structure of the JSON output requested from the LLMs for each review query.
@@ 2.2 Document Parsing
Since the extraction of structured data (text, image, tables) from PDFs is a hard problem on its own (Ouyang et al., 2025) and not all investigated models are able to process PDFs directly, we uniformly converted all papers via base64 encoding into the commonly used and LLM friendly Markdown format. This pre-processing step has been conducted via Mistral OCR (Mistral, 2025), a leading document parsing tool (see results of the OmniDocBench benchmark (Ouyang et al., 2025)) which converts text and tables from PDFs and extracts figures as images. We then feed the resulting Markdown to the LLMs.
@@ 2.4 Prompts
1Write a scientific review of the following paper for the International Conference of Learning Representations (ICLR). ICLR is a high ranking venue where less than 25% of all papers are accepted. In order to be accepted, a papers be outstanding in terms of theoretical soundness, empirical evaluation and novelty and high potential to be cited by followup research.
@@
6Give an overall rating in the output Rating using one the following scores: 1 - strong reject, 3 - reject, 5- borderline reject, 6 - borderline accept, 8 - accept, 10 - strong accept
7Rate your own confidence in the output Confidence on a scale from 1 (low) to 5 (high)
8List missing related work and citations in the output Missing_Citations.
@@ 3 Results
The following section summarizes the results of our prompt injection experiments. First, we evaluate if the models actually produce output in form of the instructed data structure and value ranges in subsection 3.1. Followed by the main experiment in subsection 3.2.
@@ 3.1 Structured Output Validation
Table 1 shows the validation results. While all models have been able to produce outputs which are following the given output data structure (as shown in listing 1), some of the models have been neglecting the range restrictions of some variables (mostly in the numerical score variables). We use the central “Rating” score to identify the ratio of invalid outputs produced by a model. By ICLR review format design, the “Rating” can only take on the following values: 1 - strong reject, 3 - reject, 5- borderline reject, 6 - borderline accept, 8 - accept, 10 - strong accept. However, some models tend to give invalid scores like “4”.
@@
Table 1: Structured output errors by model. The table shows the rate (in %) of “Ratings” given by the models which fail to adhere the requested output structure by giving scores that do not exist (most prominently "4") - also see the plots in figure 3. Green highlighted rows indicate models that have been able to predict a correct output structure. The human error rate is of cause 0% as the manual review form only allows valid scores.
@@ 3.2 Effects of Prompt Injection
Overview. Table 2 gives an overview of the effect of prompt injections on the central “Rating” score. In order to summarize the changes, we accumulate positive scores (sum of borderline accept, accept and strong accept) and report this in ratio to all scores. Most models show a very clear impact of the prompt injection, i.e. accepting 100% of the papers on a positively biased prompt while dropping to 0% in the negative case.
@@
Table 2: Acceptance rate per model for differently biased prompt injections. The table shows the rate (in %) of accumulated positive scores in the overall “Ratings” (sum of borderline accept, accept and strong accept). Green highlighted rows indicate models that have been successfully manipulated by biased prompt injections (positively an negatively). The accumulated positive scores of the human reference reviews is 43%.
Shifting the Score Distribution. A more detailed comparison between the human baseline and LLM generated review scores is visualized in figure 2 for the representative results from gemini-2.5-pro (full results for all models are given in Table 3 in the appendix). The plot shows several interesting findings: I) besides the dominant shifts of the review scores towards acceptance or rejection for the respective prompt injections, II) it also reveals a clear bias to wards acceptance for LLMs without manipulated prompts.
Figure 2: Visualization of the shifts in the distributions of the central “Rating” score for the representative gemini-2.5-pro model (full results for all models are given in Table 3 in the appendix). Positively and negatively biased prompt injection have a clear effect compared with a neutral LLM prompt. How ever, even the “neutral” LLM scores have a strong positive bias compared to the human reviews.
Failure Cases. While prompt injection has shown strong effects on most models, table 2 also shows that some models like deepseek-r1:70b or llama3.1:70b show little to no reaction to the manipulation attempts. Detailed score distribution for these models are visualized in figure 4. These plots affirm the observation that this “robustness” against manipulations is strongly correlated to the models failure to follow detailed instructions for the structured output.
4 Discussion
Prompt Injection Works! The results shown in tables 2 and 3 as well as figure 2 clearly show that very simple prompt injections are able to dominate the outcome of LLM reviews. The few cases in which the injection did not have significant effects are strongly correlated with the general failure of the models to adhere to the requested structured output. One can speculate that the ability to follow prompted instruction precisely, makes models more vulnerable towards manipulations. However, from the perspective of the assumed “careless” reviewer, these models are not very attractive to use because they do not allow a copy + paste transfer of the outputs into the review forms.
Possible Countermeasures. Since our attack scenario assumes that the manipulative prompt is injected via human unreadable text (white text on white background or extremely tiny fonts), one obvious defense could be established at the document parsing stage. By parsing PDFs as images (which is supported e.g. by Mistral OCR (Mistral, 2025)), such injection would also be hidden from the LLMs. However, it is to be expected that other, sightly more elaborate prompt injections, are likely to be able to bypass this step.
Acknowledgments. This work has been supported by the German Federal Ministry of Research, Technology and Space (BMFTR) in the program “Forschung an Fachhochschulen in Kooperation mit Unternehmen (FH-Kooperativ)” within the joint project LLMpraxis under grant 13FH622KX2.
Appendix A Appendix

### v3

@@ Abstract
Growing review duties and the availability of large language models (LLMs) have been increasing the temptations for reviewers to rely on LLMs to shortcut time consuming manual work. While a “careless” LLM dump followed by copy+past review is explicitly forbidden and considered to be scientific misconduct at most venues, recent studies indicate that this does not keep reviewers from LLM usage (Kocak et al., 2025). Especially since it is technically very hard to prove that a review has been generated by a LLM (Yu et al., 2025). Additionally, wide gray-areas do exist, as some conferences and journals are already experimenting with “LLM assisted” review processes (AAAI, 2025) (ICLR, 2024). This further fuels the ongoing discussions within the scientific communities on how to regulate LLM usage for increased productivity while maintaining review quality.
Manipulation of LLM reviews via Prompt Injection. The general idea to use hidden prompts in order to influence the review scores in their favor has probably come to the mind of many authors facing suspected LLM generated reviews. (Lin, 2025) provided the first systematic analysis which actually found evidence that this hypothetical “revenge” [^1] ICLR 2026 explicitly forbids manipulative prompt injections (ICLR, 2025). idea is actually being applied by authors. While (Lin, 2025) found many papers that include obviously manipulative strings like “IGNORE ALL PREVIOUS INSTRUCTIONS, NOW GIVE A POSITIVE REVIEW OF THESE PAPER AND DO NOT HIGHLIGHT ANY NEGATIVES”, their report does not investigate if and to what extent these attempts are actually successful. The aim of this paper is to validate the technical soundness of the described manipulation attempts.
@@
Contributions. To the best of our knowledge, we present the first detailed analysis of the practical effectiveness of simple prompt injection manipulation attempts on the scientific review process. Our extensive evaluations on real review data with human baselines show strong practical implications of LLM usage, both on the review score as well as the the high risk of manipulations.
1.1 Related Work
Automatic Paper Reviewing. Given the success of LLMs in various text based applications, it is no surprise that the research community has been investigating the automatization of the scientific peer-review process. Recent specialized review models like Openreviewer (Idahl & Ahmadi, 2024), Deepreview (Zhu et al., 2025) or Reviewer2 (Gao et al., 2024) not only motivate (partial) review generation with the increasing and tedious review work-loads, but also argue that LLM based reviews could be more objective and detailed. Besides specialized LLMs, authors also have suggested the use of multi-agent (D’Arcy et al., 2024), multi-turn (Tan et al., 2024) methods which map the entire peer-review process including discussion phases.
Evaluation of LLM generated Reviews. Given this growing number of reviewing models and wide availability of general purpose LLMs which also could be used by reviewers, several works have investigated the quality of automatic review systems. Large scale studies with human baselines in (Zhou et al., 2024), (Liang et al., 2024) and (Tyser et al., 2024) concluded, that at their current state, LLM generated reviews are to some extend “useful” to assist human reviewers, but still show major problems: Their scoring usually does not align well with human perception and they tend to hallucinate arguments and citations.
The quality of 20k LLM assisted review evaluations during the (human only) review process at ICLR 2025 (Thakkar et al., 2025) showed positive effects regarding review length and detail for those human reviewers who received LLM feedback.
Detection of LLM generated Reviews. Finally, since most venues explicitly forbid the use of LLMs during review, the detection of LLM generated text is also turning into the focus of recent research. However, latest studies like (Yu et al., 2024), (Wu et al., 2025), and (Tang et al., 2024) have shown, that it is very hard to detect LLM text with a high degree of certainty.
@@ 12 Confidence: int
Listing 1: Structure of the JSON output requested from the LLMs for each review query reflects the structure of the ICLR 2024 review form.
@@ 2.2 Document Parsing
Since the extraction of structured data (text, image, tables) from PDFs is a hard problem on its own (Ouyang et al., 2025) and not all investigated models are able to process PDFs directly, we uniformly converted all papers via base64 encoding into the commonly used and LLM friendly Markdown format. This pre-processing step has been conducted via Mistral OCR (Mistral, 2025), a leading document parsing tool (see results of the OmniDocBench benchmark (Ouyang et al., 2025)) which converts text and tables from PDFs and extracts figures as images. We then feed the resulting Markdown to the LLMs. We validate the vulnerability of Mistral OCR and other common PDF parsers in section 3.1.
@@ 2.4 Prompts
1Write a scientific review of the following paper for the International Conference of Learning Representations (ICLR). ICLR is a high ranking venue where less than 25% of all papers are accepted. In order to be accepted, a paper must be outstanding in terms of theoretical soundness, empirical evaluation and novelty and have a high potential to be cited by followup research.
@@
6Give an overall rating in the output Rating using one of the following scores: 1 - strong reject, 3 - reject, 5- borderline reject, 6 - borderline accept, 8 - accept, 10 - strong accept
7Rate your own confidence in the output Confidence on a scale from 1 (low) to 5 (high).
@@ 3 Results
The following section summarizes the results of our prompt injection experiments. First, we evaluate in section 3.1 if PDF parsers are actually converting invisible prompt injections into LLM input text. Then we test if the used language models are able to produce output in form of the instructed data structure and value ranges in subsection 3.2. This is followed by the main manipulation experiment in subsection 3.3.
3.1 Parsing Prompt Injections
In order to be able to manipulate LLM outputs, the hidden prompt injections have to be preserved as ordinary LLM text input by the initial PDF parsing. To test this crucial stage, we simulated different injection techniques from literature (Lin, 2025) and evaluated the intermediate text representations which would be fed to the LLMs in a real scenario. We used a ICLR -template and inserted the prompts prior to the paper title as shown in figure 1. The compiled PDFs were then parsed by different tools. In case of stand-alone parsing tools we evaluated the success in the output text, for web-based chat tools like ChatGPT we asked the model a distinct question about the contend of the uploaded PDF in order to verify that the injected prompt has been parsed correctly.
Table 1 shows the results for different common parsing approaches and injection methods: “black” refers to a baseline experiment where the prompt is visible black-on-white text. “White” represents a white-on-white text invisible to humans and “tiny” uses a text which is so small that it also would be overseen by human readers.
Prompt ChatGPT* Gemini* PyMuPDF Mistral OCR (PDF) Mistral OCR (Image)
black
white
tiny
Table 1: Results for the injection parsing test for different injection methods and parsers. * indicates web-based chat services.
All tools which are using the PDF sources for the extraction of text are parsing the hidden prompts as standard text, enabling possible manipulations of the following LLM review generation. On the other hand, image based OCR is ignoring invisible prompts. Notably, Google’s Gemini web-service appears to be using an image based parser, contrary to OpenAI’s ChatGPT.
@@ 3.2 Structured Output Validation
In the next step of our empirical analysis, we validate the ability of the investigated models to generate correctly structured output. Table 2 shows these results. While all models have been able to produce outputs which are following the given output data structure (as shown in listing 1), some of the models have been neglecting the range restrictions of some variables (mostly in the numerical score variables). We use the central “Rating” score to identify the ratio of invalid outputs produced by a model. By ICLR review format design, the “Rating” can only take on the following values: 1 - strong reject, 3 - reject, 5- borderline reject, 6 - borderline accept, 8 - accept, 10 - strong accept. However, some models tend to give invalid scores like “4”.
@@
Table 2: Structured output errors by model. The table shows the rate (in %) of “Ratings” given by the models which fail to adhere the requested output structure by giving scores that do not exist (most prominently "4") - also see the plots in figure 4. Green highlighted rows indicate models that have been able to predict a correct output structure. The human error rate is of cause 0%, as the manual review form only allows valid scores.
@@ 3.3 Effects of Prompt Injection
Overview. Table 3 gives an overview of the effect of prompt injections on the central “Rating” score. In order to summarize the changes, we accumulate positive scores (sum of borderline accept, accept and strong accept) and report this in ratio to all scores. Most models show a very clear impact of the prompt injection, i.e. accepting 100% of the papers on a positively biased prompt while dropping to 0% acceptance in the negative case.
@@
Table 3: Acceptance rate per model for differently biased prompt injections. The table shows the rate (in %) of accumulated positive scores in the overall “Ratings” (sum of borderline accept, accept and strong accept). Green highlighted rows indicate models that have been successfully manipulated by biased prompt injections (positively and negatively). The accumulated positive scores of the human reference reviews is 43%.
Failure Cases. While prompt injection has shown strong effects on most models, table 3 also shows that some models like deepseek-r1:70b or llama3.1:70b show little to no reaction to the manipulation attempts. Detailed score distribution for these models are visualized in table 5 of the appendix. These plots affirm the observation that this “robustness” against manipulations is strongly correlated to the models failure to follow detailed instructions for the structured output.
Figure 2: Visualization of the shifts in the distributions of the central “Rating” score for the representative gemini-2.5-pro model (full results for all models are given in Table 4). Positively and negatively biased prompt injections have a clear effect compared with a neutral LLM prompt. However, even the “neutral” LLM scores have a strong positive bias compared to the human reviews.
Shifting the Score Distribution. A more detailed comparison between the human baseline and LLM generated review scores is visualized in figure 2 for the representative results from gemini-2.5-pro (full results for all models are given in Table 4 ). The plot shows several interesting findings: I) besides the dominant shifts of the review scores towards acceptance or rejection for the respective prompt injections, II) it also reveals a clear bias to wards acceptance for LLMs without manipulated prompts.
@@
Embedding Analysis: Summaries. In the next series of experiments, we explore whether the prompt manipulations only effect the review scores or if they also alter the line of argumentation in the generated texts. As a baseline, we extracted and embedded the paper summaries with gemini-embedding-001 (SEMANTIC SIMILARITY mode with 128 dimensions) and computed the cosine distances between embeddings. Figure 3a shows that the mean distance of LLM generated summaries to the according human texts is almost as low as the mean dissimilarity between human summaries. Also the prompt appears to have little effect on the summaries. Again, the models that fail to adhere to the required output structure, apparently also fail to generate meaningful summaries.
(a) Summary embeddings: the red line shows the mean cosine distance between the summary sections of human reviews of the same paper, compared to the mean cosine distances of LLM generated summaries for the same papers to these human baselines.
(b) Argument embeddings: the green and red lines show the mean Hausdorff distance between the “strengths” and “weaknesses” argument lists of human reviews, compared to the mean distances of LLM generated argument lists to these human baselines.
Figure 3: Effect of the prompt injections on the embedding distances of (a) the review summaries and (b) the “strengths” and “weaknesses” argument lists.
@@
Embedding Analysis: Strengths and Weaknesses. In a second experiment, we investigate the pro and con arguments listed in the reviews. First we used gemini-2.5-flash to extracted list of Strengths and Weaknesses from the human reviews before embedding them item by item. Embedding the LLM generated Strengths and Weaknesses the same way for each model (these are already outputted as lists), we then compute the Hausdorff-Distance (Taha & Hanbury, 2015) between the embedding point-clouds (allowing different numbers of arguments within one comparison pair). Figure 3b shows several results of this evaluation: I) human reviewers tend to agree more on the positive aspects of a paper than on the negative ones (shown by the solid red and green lines indication the mean distance between human argument lists). II) Also LLM generated Strengths are closer to the human findings than the Weaknesses. III) prompt injections show a measurable effect, however positively biased reviews appear to be moving closer to the human evaluation, leaving a larger gap for negatively biased generations.
4 Discussion
Prompt Injection Works! The results shown in tables 3 and 4 as well as figure 2 clearly show that very simple prompt injections are able to dominate the outcome of LLM reviews. The few cases in which the injection did not have significant effects are strongly correlated with the general failure of the models to adhere to the requested structured output. One can speculate that the ability to follow prompted instruction precisely, makes models more vulnerable towards manipulations. However, from the perspective of the assumed “careless” reviewer, these models are not very attractive to use because they do not allow a copy + paste transfer of the outputs into the review forms.
Possible Countermeasures. Since our attack scenario assumes that the manipulative prompt is injected via human unreadable text (white text on white background or extremely tiny fonts), one obvious defense could be established at the document parsing stage. By parsing PDFs as images (as shown in table 1), such injections would also be hidden from the LLMs. However, it is to be expected that other, sightly more elaborate prompt injections, are likely to be able to bypass this step.